OpenStudy (anonymous):

Given the following two points on a line, write the equation of the line in slope-intercept and point-slope form. (1, 4) and (-3, -2)

OpenStudy (texaschic101):

we need to find the slope using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (1,4) x1 = 1 and y1 = 4 (-3,-2) x2 = -3 and y2 = -2 now we sub slope(m) = (-2-4) / (-3-1) slope(m) = -6/-4 which equals 3/2 point-slope form : y - y1 = m(x - x1) slope(m) = 3/2 you can choose either of your set of points (1,4) x1 = 1 and y1 = 4 now we sub y - 4 = 3/2(x - 1) ====>POINT-SLOPE FORM y - 4 = 3/2(x - 1) -- distribute through the parenthesis y - 4 = 3/2x - 3/2 y = 3/2x - 3/2 + 4 y = 3/2x - 3/2 + 8/2 y = 3/2x + 5/2 =====> SLOPE-INTERCEPT FORM

OpenStudy (texaschic101):

do you have any questions at all ?